# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

### 30 Jan 2024 v2

@@ 3.4. RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
Figure 7. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) Should that work and its derivatives be considered yours? and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 7. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_B (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) Should that work and its derivatives be considered yours? and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
RQ4_C: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the AI model creators?
@@
Figure 8. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the AI model creators?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 8. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_C (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the AI model creators?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
@@
According to our survey results, only 26.80% of the participants believed that the work and its derivatives should be considered the property of the creators of AI models. To examine the relationship between the answers to RQ4_C on the Likert scale and other secondary variables, we conducted an ordinal logistic regression (OLR), as illustrated in Figure 8.
Our analysis suggests a strong association (with a p-value below 0.05, presented in Table 4) between agreement on AI companies owning the resulting artwork and art practice, race, gender, familiarity with AI art models, and compensation.
@@ 4. Discussion
In this study, we surveyed how artists feel about Generative Artificial Intelligence (AI) in art. Most participants believe that AI model creators should be required to disclose the art and images used to train them – an aspect of Generative AI tools that is currently far from transparent (see Sec. 2.4). However, only a minority of participants see AI art models as a positive development in art, while the majority see them as a threat to art workers.
@@
Regarding compensation for using their artwork to train AI art models, over half of the participants did not require any compensation but were concerned about who was profiting from their artwork. Tentatively, we think this answer could reflect a belief within the artistic community: that art is the joint responsibility of all people (Huyghe, 1970). Corollary to that, there would be some onus on artists to share their work with each other and with the public, and there may be resistance to perceived exploitation of that practice, especially in the service of profit. We think this interpretation is consistent with the result that some participants stated that they do not require profit but do not want for-profit companies to profit from their art. Others considered some compensation fair, while a minority felt taxing companies and individuals who profit from AI art outputs would be a fair option. Finally, some participants stated they would feel uncomfortable with the compensation options listed. Understanding how the field of Generative AI falls short in respecting the values of the people who make the cultural artifacts it relies upon – and how that could be changed – is an important ethical question that we hope to expand on in future work.
@@
In our future work, we plan to increase the sample size and expand it to a more representative sample. This will increase the statistical power of our study and also allow us to look at how opinions change over time. Moreover, we would like to put the current flourishing of Generative AI within the context of art history and taxonomy. We believe that a longer timeline would help consolidate, unify, and clarify the relationship between technology and art, which we only briefly touched upon in Section 2.5. While it is essential to understand what artists think about Generative AI, we need to collaborate with experts from various fields, including art, history, philosophy, psychology, cognitive science, and others, to understand the implications of Generative AI on our understanding of art and its role in our lives.
Acknowledgements.
@@
Institutional Review Board Approval: The survey in this project is CHRBSS (Behavioral) STUDY00002379, approved by the University of Vermont I.R.B. on 3/14/2023. The authors would like to thank Maria Sckolnick, John Meluso, Kendall Fortney, Randall Harp, Laurent Hébert-Dufresne, Jean-Gabiel Young, and Winn Gillen, for their comments on early versions of this work. The Alfred P. Sloan Foundation, The National Science Foundation award #2242829, and MassMutual under the MassMutual Center of Excellence in Complex Systems and Data Science support this work. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the aforementioned financial supporters.
Appendix A Research Methods
A.1. Survey Methodology
Our anonymous original survey of artists is an observational study (N=516) of artists over the age of 18 years. To compare, we present a table of the distributions of the US Census art workforce 5. Incomplete surveys were dropped from the analysis, leaving a final N of 459.
Survey participants start by consenting to the IRB-approved survey information sheet and consent agreement. We then asked participants to answer several questions about their artistic identity and financial relationship with art. Before continuing, participants are asked to read a brief describing AI art and present an example. We then ask participants to disclose their familiarity with AI art models and tools. Then, we asked participants several questions about who they think “owns” art generated by AI models rated on a 5-point Likert scale. We ask several questions that invite artists to identify their ideal compensation criteria for their art being included in AI art models. We then ask participants to assess if AI art models are positive for their field, if they are a threat to their field, and if they think model creators should be required to disclose their data sources on a 5-point Likert scale. Lastly, we collect demographic information on the survey participants’ backgrounds.
Phase 2 Sampling: Phase 2 (n = 264) was collected via a targeted survey panel conducted by Qualtrics from September 12-23, 2023 on the Qualtrics online survey platform. A link to the complete survey and survey questions is available in Section B.
Descriptive statistics about the U.S. artist workforce are categorized as artists-related workers (US Census code 27-1010), which can be seen in Table 5. Data comes from the 2020 US Census, US Census Bureau, and ACS PUMS 5-Year Estimate. The occupational categories include Art Directors, Craft Artists, Fine Artists, Painters, Sculptors, Illustrators, Special Effects Artists and Animators, Artists and Related Workers, All Other Designers, Commercial and Industrial Designers, Fashion Designers, Floral Designers, Graphic Designers, Interior Designers, Merchandise Displayers and Window Trimmers, Set and Exhibit Designers, Designers, All Other. This data demonstrates the number of US workers potentially implicated by Generative AI art.
Type Number
U.S. artist workforce 144,000
Avg. Age 41
Avg. Salary \$57,958
Avg Male Salary 67,353
Avg Female Salary 45,618
Male 56.70%
Female 43.30%
White (Non-Hispanic) 73.20%
Asian (Non-Hispanic) 7.92%
Black 3.65%
Two or more 3.44%
White (Hispanic) 2.88%
American Indian 0.01%
Native Hawaiian 0.13%
Native other 0.04%
A.2. Data
Original Data
Our original survey of artists is a convenience survey (N=459). For comparison of the distributions of the US art workforce) based on the demographics is presented in Table 5.
We transformed all survey response variables of interest into numerical form to analyze our survey results (survey questions 8,9,10,13,14,15 as seen in Section B). All Likert survey questions were converted from ‘Strongly Agree,’ ‘Agree,’ ‘Neutral,’ ‘Disagree,’ and ‘Strongly Disagree’ to an ordinal scale of 1,2,3,4,5. For our ordinal logistic regression analysis, we combine ‘Strongly Agree’ and ‘Agree’ to be equal to 3, ’Neutral’ to be equal to 2, and ‘Disagree’ and ‘Strongly Disagree’ to be equal to 1. In the original logistic regression, we reduced the dimensionality of the Likert scale to investigate the positive, neutral, and negative sentiment of the response questions in relation to our secondary variables.
All nominal and categorical variables (survey questions 1,2,3,4,5,11,12,16,17,18,19 as seen in Section B) were transformed into binary variables. Survey questions that included an “other” option or write-in answers were also categorized (e.g., in question 2, we added an additional art practice, which included digital art from the write-in answers).
Demographic questions (survey questions 16, 17, 18, 19 as seen in Section B) included asking survey participants their gender identity to identify whether they were white or a person of color, choosing their age range and country of residence. All demographic questions were transformed into binary variables.
A.3. Analytical Methods
Kruskal–Wallis Test The Kruskal-Wallis H-test is a non-parametric test that extends the Mann-Whitney U-test to compare differences between groups for an ordinal or continuous dependent variable. The explanatory groups may be ranked or unranked (Statistics, 2015a). Its null hypothesis is stochastic homogeneity, which is equivalent to equality of the expected value of rank means, conditional on the homogeneity of rank variances (Vargha and Delaney, 1998).
Odds ratio In general, the odds of an event are the probability of success over the probability of failure. This can be written
(1) $\frac{p}{1-p}$
Ordinal Logistic Regression Ordinal logistic regression (OLR) is a generalization of regression (either multiple linear or binomial logistic) used for an ordered categorical response variable and any number of categorical or numeric secondary variables (Statistics, 2015b). OLR finds the logit (also known as log-odds) of the dependent variable (Y) being less than or equal to a certain level (denoted j) (Bruin, 2011b). which can be written
(2) $logit(P(Y\leq j))=log\frac{P(Y\leq j)}{P(Y>j)}$
It is important to note that OLR assumes proportional odds, meaning the effect of the secondary variable(s) is the same between all levels of the dependent variable (Unit, 2016).
Appendix B Survey Questions
@@
    Please read the following description of Artificial Intelligence (AI) art before continuing: AI art is any computer-generated artwork created through the use of artificial intelligence. There are several ways to generate AI art. Models used to generate AI art: Artificial intelligence algorithms use two or more images to train (training is a process used where an algorithm is fed data to learn from) a model to generate new AI art (e.g., Lensa, DeepDream). Large language models use artificial intelligence, which uses large amounts of text data to train a model to generate new AI language (e.g., GPT-3). Text-to-image models use textual prompts (e.g., “a yellow elephant”) from a user to generate AI art that depicts the textual prompt. The AI art is trained on large amounts of image and text data to train the model and generate new AI art (e.g., DALL-E, Imagen). Example: Generated by DALL-E mini/Craiyon using the prompt “a cat in the style of Van Gogh.” According to its creators, “[t]he model was trained on unfiltered data from the Internet, limited to pictures with English descriptions.” Includes image seen in Figure 1.
@@
Appendix C Full table of descriptive statistics
A complete table of descriptive statistics can be found in Table 6.
@@ Appendix C Full table of descriptive statistics
Appendix D All tables for ordinal logistic regressions
We ran an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to each primary research question). Our secondary variables, the results of these original logistic regressions, their odds-ratios, their p-values, and their confidence intervals (CI) can be seen in Tables 7, 8, 9, 10, 11, 12.
@@ Appendix D All tables for ordinal logistic regressions
Artist 0.364481 0.913316 0.576963 ¡0.001
Purchased art 0.331723 1.004549 0.577263 ¡0.001
Person of color 1.23678 3.085823 1.953582 ¡0.001
Professional artist 0.840948 2.063429 1.317284 ¡0.001
Used AI art models 3.4805 7.703421 5.178007 ¡0.001
@@
Female 0.417108 0.868241 0.601789 ¡0.001
Non-binary 0.420314 2.076117 0.934141 ¡0.001
Age 50-64 0.5506 1.369427 0.868336 ¡0.001
Age 65 plus 0.346227 1.0817 0.611975 ¡0.001
@@
None, donate use of artwork to train models 4.162129 57.308125 15.444217 ¡0.001
@@
Artist 0.673062 1.734431 1.080453 ¡0.001
Purchased art 0.56515 1.850283 1.022589 ¡0.001
Person of color 0.835831 2.157276 1.342802 ¡0.001
Professional artist 0.511389 1.27991 0.809031 ¡0.001
Used AI art models 0.381466 0.843053 0.567094 ¡0.001
Male 0.68315 1.452222 0.996035 ¡0.001
Female 0.678301 1.484131 1.003338 ¡0.001
Non-binary 0.41613 2.447633 1.009224 ¡0.001
Age 50-64 0.399295 1.029979 0.641299 ¡0.001
Age 65 plus 0.385716 1.286586 0.704455 ¡0.001
Age 18-29 1.337791 4.055724 2.329316 ¡0.001
Age 30-49 0.829556 1.938754 1.26819 ¡0.001
@@
Artist 0.618355 1.949844 1.098041 ¡0.001
Purchased art 0.851732 3.164821 1.641822 ¡0.001
Person of color 0.809301 2.740342 1.489215 ¡0.001
Professional artist 0.640443 2.00734 1.133837 ¡0.001
Used AI art models 0.38645 1.045068 0.635505 ¡0.001
Male 0.587018 1.506886 0.940515 ¡0.001
Female 0.669859 1.797992 1.097452 ¡0.001
Non-binary 0.30543 2.406881 0.857399 ¡0.001
Age 50-64 0.236998 0.712557 0.410944 ¡0.001
Age 65 plus 0.325902 1.57366 0.716142 ¡0.001
Age 18-29 1.427261 5.790101 2.874715 ¡0.001
Age 30-49 1.061875 2.97578 1.777612 ¡0.001
Familiarity AI art gen 1.363085 5.05368 2.624614 ¡0.001
Familiarity AI/DL 0.879414 6.605245 2.410134 ¡0.001
Familiarity text to image 0.131881 1.59315 0.458373 ¡0.001
@@
Female 0.540109 1.114689 0.775921 ¡0.001
Non-binary 0.304406 1.426784 0.659031 ¡0.001
Age 50-64 0.433393 1.072792 0.681865 ¡0.001
Age 65 plus 0.715666 2.230618 1.263478 ¡0.001
@@
Artist 0.414702 1.019737 0.650297 ¡0.001
Purchased art 0.327522 0.974205 0.564866 ¡0.001
Person of color 1.120118 2.725496 1.747248 ¡0.001
Professional artist 0.702605 1.699884 1.092862 ¡0.001
Used AI art models 2.034051 4.377637 2.984013 ¡0.001
Male 1.41509 2.883435 2.01998 ¡0.001
Female 0.314178 0.66174 0.455965 ¡0.001
Non-binary 0.389118 1.943853 0.869706 ¡0.001
@@
Appendix E Contingency Tables:
All contingency tables for dependent variables with secondary variables can be seen in Tables 13, 14, 15, 16, 18, 19, 20, 21, 17, 22, 23, 24, and 25.
@@ Appendix E Contingency Tables:
Appendix F Tests for Independence
F.0.1. Kruskal-Wallis test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are categorical, we ran Kruskall-Wallis tests for independence. The results of these tests can be seen in Tables 26, 27, 30, 31, 28, 29.
@@ F.0.1. Kruskal-Wallis test for independence:
F.0.2. Cochran-Armitage test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are binary, we ran Cochran-Armitage tests for independence. The results of these tests can be seen in Tables 32, 33, 36, 37, 34, and 35.

### 14 May 2024 v4

@@ 4. Results
Figure 6. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4B (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh) Should that work and its derivatives be considered yours? and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
@@
Figure 7. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4C (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the AI model creators?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
Finally, we look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. The analysis for RQ4 is related to survey questions No. 8, 9, and 10 (all survey questions can be seen in Appendix I).
RQ4A: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?
Here, we explore the scenario where the owner of the Generative AI artwork should be the person who used the AI model to generate the artwork. We see that 39.22% of participants agreed that the work and its derivatives should be considered the property of the person who used the AI model to generate the artwork. We conducted an ordinal logistic regression (OLR) to investigate the relationship between our dependent variable (Likert scale responses to RQ4A) and a subset of our secondary variables, represented by the odds ratio depicted in Figure 5.
We see evidence of a strong association (p-value $<$ 0.05) between agreeing and our secondary variables. Namely, males, persons of color, and participants who have used Generative AI to make art were more likely to agree that the owner of the artwork in this scenario should be the person who used the AI model to generate the artwork. Whereas females, participants who had purchased a piece of artwork, and participants who said that they were not comfortable with any type of compensation for the use of their artwork to train a generative AI model were less likely to agree that the owner of the artwork in this scenario should be the person who used the AI model to generate the artwork.
RQ4B: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh), Should that work and its derivatives be considered yours?
Here, we explore the scenario where the owner of the Generative AI artwork should be the artist whose style is being recognizable and mimicked by the Generative AI art model. Our survey shows that 41.39% of the participants agreed that the work and its derivatives should be considered their property (the artist whose style is being recognizable and mimicked by the Generative AI art model).
To investigate the relationship between the Likert scale responses to RQ4B and the secondary variables, we conducted an ordinal logistic regression (OLR), presented in Figure 6 displaying the odds ratios derived from the analysis. Our results show that there is a strong relationship (p-value $<$ 0.05, all detailed odds ratios can be seen in Appendix III) between agreeing that the artist whose style is recognizable in the output should be the owner of the art and a subset of our secondary variables. We see that participants who are persons of color are more likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model. Whereas participants who have used Generative AI to make art were less likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model.
RQ4C: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the AI model creators?
@@
Finally, we look at our final scenario where the owner of the Generative AI artwork should be the AI model creators. According to our survey results, only 26.80% of the participants believed that the work and its derivatives should be considered the property of the creators of AI models. To examine the relationship between the answers to RQ4C on the Likert scale and other secondary variables, we conducted an ordinal logistic regression (OLR), as illustrated in Figure 7.
Our analysis suggests a strong relationship (p-value $<$ 0.05, all odds ratios can be seen in Appendix III) between agreement on AI companies owning the resulting artwork and a subset of our secondary variables. Participants who considered themselves tattooists (note that tattooists were a small n (n=9) with a fairly large range in the confidence interval bars represented in Figure 7), males, persons of color, or those who said they would donate their artwork to train generative AI models were more likely to agree that the owner of the Generative AI artwork should be the AI model creators. Whereas participants with a general familiarity with Generative AI Art, females, and non-binary participants were less likely to agree that the owner of the Generative AI artwork should be the AI model creators.
@@ 5. Discussion
In this study, we surveyed how artists feel about Generative Artificial Intelligence (AI) in art. Most participants believe that AI model creators should be required to disclose the art and images used to train them – an aspect of Generative AI tools that is far from transparent. However, only a minority of participants see AI art models as a positive development in art, while the majority see them as a threat to art workers.
@@
Regarding compensation for using their artwork to train AI art models, over half of the participants did not require any compensation but were concerned about who was profiting from their artwork. One concern is that artists will be reluctant to share their work with each other and with the public going forward, and there may be resistance to perceived exploitation of that practice, especially in the service of profit. We think this interpretation is consistent with the result that some participants stated that they do not require profit but do not want for-profit companies to profit from their art. Others considered some compensation fair, while a minority felt taxing companies and individuals who profit from AI art outputs would be a fair option. Finally, some participants stated they would feel uncomfortable with the compensation options listed. Understanding how the field of Generative AI falls short in respecting the values of the people who make the cultural artifacts it relies upon – and how that could be changed – is an important ethical question that we hope to expand on in future work.
@@
In our future work, we plan to increase the sample size and expand it to a more representative sample. This will increase the statistical power of our study and also allow us to look at how opinions change over time. Moreover, we would like to put the current flourishing of Generative AI within the context of art history and taxonomy. We believe that a longer timeline would help consolidate, unify, and clarify the relationship between technology and art. While it is essential to understand what artists think about Generative AI, we need to collaborate with experts from various fields, including art, history, philosophy, psychology, cognitive science, and others, to understand the implications of Generative AI on our understanding of art and its role in our lives.
6. Acknowledgements
@@
We are thankful to the anonymous reviewers for their insightful comments. Institutional Review Board Approval: The survey in this project is CHRBSS (Behavioral) STUDY00002379, approved by the University of Vermont I.R.B. on 3/14/2023. The authors would like to thank Maria Sckolnick, John Meluso, Kendall Fortney, Randall Harp, Laurent Hébert-Dufresne, Jean-Gabiel Young, and Winn Gillen, for their comments on early versions of this work. The Alfred P. Sloan Foundation, The National Science Foundation award #2242829, and MassMutual under the MassMutual Center of Excellence in Complex Systems and Data Science support this work. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the aforementioned financial supporters.
Appendix A Appendix I: Survey Questions
@@
    Please read the following description of Artificial Intelligence (AI) art before continuing: AI art is any computer-generated artwork created through the use of artificial intelligence. There are several ways to generate AI art. Models used to generate AI art: Artificial intelligence algorithms use two or more images to train (training is a process used where an algorithm is fed data to learn from) a model to generate new AI art (e.g., Lensa, DeepDream). Large language models use artificial intelligence, which uses large amounts of text data to train a model to generate new AI language (e.g., GPT-3). Text-to-image models use textual prompts (e.g., “a yellow elephant”) from a user to generate AI art that depicts the textual prompt. The AI art is trained on large amounts of image and text data to train the model and generate new AI art (e.g., DALL-E, Imagen). Example: Generated by DALL-E mini/Craiyon using the prompt “a cat in the style of Van Gogh.” According to its creators, “[t]he model was trained on unfiltered data from the Internet, limited to pictures with English descriptions.”
@@
Appendix B Appendix II: Full table of descriptive statistics
A complete table of descriptive statistics for our participants in the survey can be found in Table 5.
Descriptive statistics about the US artist workforce who are categorized as artists-related workers (US Census code 27-1010) can be seen in Table 4. Data comes from the 2020 US Census, US Census Bureau, and ACS PUMS 5-Year Estimate. Descriptive statistics about the U.S. artist workforce are categorized as artists-related workers (US Census code 27-1010), which can be seen in Table 4. Data comes from the 2020 US Census, US Census Bureau, and ACS PUMS 5-Year Estimate. The occupational categories include Art Directors, Craft Artists, Fine Artists, Painters, Sculptors, Illustrators, Special Effects Artists and Animators, Artists and Related Workers, All Other Designers, Commercial and Industrial Designers, Fashion Designers, Floral Designers, Graphic Designers, Interior Designers, Merchandise Displayers and Window Trimmers, Set and Exhibit Designers, Designers, All Other. This data demonstrates the number of US workers potentially implicated by Generative AI art.
Appendix C Appendix III: All tables for ordinal logistic regressions
We ran an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to each primary research question). Our secondary variables, the results of these original logistic regressions, their odds-ratios, their p-values, and their confidence intervals (CI) can be seen in Tables 6, 7, 8, 9, 10, 11.
Appendix D Appendix IV: Contingency Tables
All contingency tables for dependent variables with secondary variables can be seen in Tables 12, 13, 14, 15, 17, 18, 19, 20, 16, 21, 22, 23, and 24.
Appendix E Appendix V: Tests for Independence
E.0.1. Kruskal-Wallis test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are categorical, we ran Kruskall-Wallis tests for independence. The results of these tests can be seen in Tables 25, 26, 29, 30, 27, 28.
Kruskal–Wallis Test The Kruskal-Wallis H-test is a non-parametric test that extends the Mann-Whitney U-test to compare differences between groups for an ordinal or continuous dependent variable. The explanatory groups may be ranked or unranked (website, 2018). Its null hypothesis is stochastic homogeneity, which is equivalent to equality of the expected value of rank means, conditional on the homogeneity of rank variances (Vargha and Delaney, 1998).
E.0.2. Cochran-Armitage test for independence:
To test for independence between dependent variables that were ordinal with variables of interest that are binary, we ran Cochran-Armitage tests for independence. The results of these tests can be seen in Tables 31, 32, 35, 36, 33, and 34.
Type Number
U.S. artist workforce 144,000
Avg. Age 41
Avg. Salary \$57,958
Avg Male Salary 67,353
Avg Female Salary 45,618
Male 56.70%
Female 43.30%
White (Non-Hispanic) 73.20%
Asian (Non-Hispanic) 7.92%
Black 3.65%
Two or more 3.44%
White (Hispanic) 2.88%
American Indian 0.01%
Native Hawaiian 0.13%
Native other 0.04%
@@ E.0.2. Cochran-Armitage test for independence:
Artist 0.364481 0.913316 0.576963 $<$0.001
Purchased art 0.331723 1.004549 0.577263 $<$0.001
Person of color 1.23678 3.085823 1.953582 $<$0.001
Professional artist 0.840948 2.063429 1.317284 $<$0.001
Used AI art models 3.4805 7.703421 5.178007 $<$0.001
@@
Female 0.417108 0.868241 0.601789 $<$0.001
Non-binary 0.420314 2.076117 0.934141 $<$0.001
Age 50-64 0.5506 1.369427 0.868336 $<$0.001
Age 65 plus 0.346227 1.0817 0.611975 $<$0.001
@@
None, donate use of artwork to train models 4.162129 57.308125 15.444217 $<$0.001
@@
Artist 0.673062 1.734431 1.080453 $<$0.001
Purchased art 0.56515 1.850283 1.022589 $<$0.001
Person of color 0.835831 2.157276 1.342802 $<$0.001
Professional artist 0.511389 1.27991 0.809031 $<$0.001
Used AI art models 0.381466 0.843053 0.567094 $<$0.001
Male 0.68315 1.452222 0.996035 $<$0.001
Female 0.678301 1.484131 1.003338 $<$0.001
Non-binary 0.41613 2.447633 1.009224 $<$0.001
Age 50-64 0.399295 1.029979 0.641299 $<$0.001
Age 65 plus 0.385716 1.286586 0.704455 $<$0.001
Age 18-29 1.337791 4.055724 2.329316 $<$0.001
Age 30-49 0.829556 1.938754 1.26819 $<$0.001
@@
Artist 0.618355 1.949844 1.098041 $<$0.001
Purchased art 0.851732 3.164821 1.641822 $<$0.001
Person of color 0.809301 2.740342 1.489215 $<$0.001
Professional artist 0.640443 2.00734 1.133837 $<$0.001
Used AI art models 0.38645 1.045068 0.635505 $<$0.001
Male 0.587018 1.506886 0.940515 $<$0.001
Female 0.669859 1.797992 1.097452 $<$0.001
Non-binary 0.30543 2.406881 0.857399 $<$0.001
Age 50-64 0.236998 0.712557 0.410944 $<$0.001
Age 65 plus 0.325902 1.57366 0.716142 $<$0.001
Age 18-29 1.427261 5.790101 2.874715 $<$0.001
Age 30-49 1.061875 2.97578 1.777612 $<$0.001
Familiarity AI art gen 1.363085 5.05368 2.624614 $<$0.001
Familiarity AI/DL 0.879414 6.605245 2.410134 $<$0.001
Familiarity text to image 0.131881 1.59315 0.458373 $<$0.001
@@
Female 0.540109 1.114689 0.775921 $<$0.001
Non-binary 0.304406 1.426784 0.659031 $<$0.001
Age 50-64 0.433393 1.072792 0.681865 $<$0.001
Age 65 plus 0.715666 2.230618 1.263478 $<$0.001
@@
Artist 0.414702 1.019737 0.650297 $<$0.001
Purchased art 0.327522 0.974205 0.564866 $<$0.001
Person of color 1.120118 2.725496 1.747248 $<$0.001
Professional artist 0.702605 1.699884 1.092862 $<$0.001
Used AI art models 2.034051 4.377637 2.984013 $<$0.001
Male 1.41509 2.883435 2.01998 $<$0.001
Female 0.314178 0.66174 0.455965 $<$0.001
Non-binary 0.389118 1.943853 0.869706 $<$0.001